# Progressive Confident Masking Attention Network for Audio-Visual Segmentation

### v1

@@ Abstract
Audio and visual signals typically occur simultaneously, and humans possess an innate ability to correlate and synchronize information from these two modalities. Recently, a challenging problem known as Audio-Visual Segmentation (AVS) has emerged, intending to produce segmentation maps for sounding objects within a scene. However, the methods proposed so far have not sufficiently integrated audio and visual information, and the computational costs have been extremely high. Additionally, the outputs of different stages have not been fully utilized. To facilitate this research, we introduce a novel Progressive Confident Masking Attention Network (PMCANet). It leverages attention mechanisms to uncover the intrinsic correlations between audio signals and visual frames. Furthermore, we design an efficient and effective cross-attention module to enhance semantic perception by selecting query tokens. This selection is determined through confidence-driven units based on the network’s multi-stage predictive outputs. Experiments demonstrate that our network outperforms other AVS methods while requiring less computational resources.
Index Terms:
audio-visual segmentation, cross-modal attention mechanism, dynamic network, multi-modality
I Introduction
@@
As mentioned earlier, the number of unlabeled videos is rapidly increasing. Research on these tasks primarily focuses on semi-supervised learning [18] and unsupervised learning, particularly self-supervised learning [19, 20, 21, 22, 23, 24]. However, due to the lack of patch-level or pixel-level annotation, namely, supervision, these tasks are limited to image-level classification, and their applications are severely constrained. For years, researchers did not pay much attention to supervised learning and pixel-level annotation. Zhou et al. [25, 26] were the first to propose the Audio-Visual Segmentation (AVS) task and Audio-Visual Segmentation with Semantic (AVSS) task, providing benchmark datasets specifically for these tasks. There is a similar task called Sound Source Localization (SSL) [20, 18, 21, 27, 22, 23, 28, 29], which aims to locate the regions in the visual scene that are relevant to the audio signals. Although it is not pixel-level, it is annotated at the patch-level. The results are typically represented by heat maps, which are obtained through similarity matrices or Class Activation Mapping (CAM) [30].
@@
II RELATED WORK
II-A Audio-Visual Correspondence
@@
II-B Sound Source Localization
@@
Figure 2: Overview of the pipeline. The network takes the video clip as input and separates it into visual frames and audio spectrogram. The outputs of the audio and visual encoders are denoted as $F_{i}$ and $A$, respectively. The features are first integrated by $AVGA$s and are further processed by $QSCA$s to generate guide features. Meanwhile, the multi-stage outputs, which are aggregated by $GF$s, will pass through $CIM$s to generate confidence masks and then be sent to $QSCA$s.
II-C Audio-Visual Attention
@@
III PROPOSED METHOD
III-A Overview Structure of PCMANet
@@
III-B Audio-Visual Group Attention
With the extracted audio and visual features, it is intuitive to assess whether there exists a correlation between the audio signal and spatial localization. Previous studies[22, 18] introduced an audio-visual attention mechanism by computing the cosine similarity between audio and visual features at the pixel level, and it was found to be effective in tasks such as sound source localization and separation. Drawing inspiration from these findings, we design a novel module named Audio-Visual Group Attention (AVGA).
@@ III-B Audio-Visual Group Attention
Figure 3: Structure of Audio-Visual Group Attention (AVGA).
III-C Query-Selected Cross-Attention
@@
Figure 4: Structure of Query-Selected Cross-Attention (QSCA).
Furthermore, the visual tokens are redundant, and the computational cost of the attention operation is exceedingly high. To address this challenge, we propose a new query-selected mechanism. Instead of using all tokens, the network selects tokens that lack confidence for segmentation [50]. The criteria for selection are measured by the Confidence-Induced Masking unit (see Section III-D). The entire process can be formulated as follows:
Figure 5: An example of the QSCA and Confidence-Induced Masking (CIM).
@@
where $N^{\prime}=rN$, the ratio $r$ gradually increases from 00 to $1$. Experiments demonstrate that in the later stages of training, $r$ exceeds $90\%$, indicating that a significant portion of the tokens is masked, resulting in substantial savings in computational resources. Figure 5 provides an example of the confidence masks and the selected query.
III-D Confidence-Induced Masking and Guided Fusion
@@
The Guided Fusion module essentially combines adjacent features, similar to FPN [51]. The key distinction lies in its utilization of the output from the QSCA module, which has already integrated audio features. This modification ensures that the prediction generation process focuses more on the sounding objects.
IV SUPERVISION
TABLE I: Quantitative comparison on AVS dataset. The best result is highlighted in red. The arrow $\uparrow$ of the metrics indicates the value is higher the result is better. ResNet-50, PVT-v2, and Swin refer to the ResNet-50 [37], PVT-v2 [39], and Swin Transformer [52], respectively.
Model S4 MS3
$F_{m}$$\uparrow$ $mIoU$$\uparrow$ $F_{m}$$\uparrow$ $mIoU$$\uparrow$
ResNet-50 LVS .510 37.94 .330 29.45
3DC .759 57.10 .503 36.92
iGAN .778 61.59 .544 42.89
AVS .848 72.79 .578 47.88
PCMA .858 75.24 .648 54.77
PVT-v2 / Swin MSSL .663 44.89 .363 26.13
SST .801 66.29 .572 42.57
LGVT .873 74.94 .593 40.71
AVS .879 78.74 .645 54.00
PCMA .893 80.10 .708 60.35
TABLE II: Quantitative comparison on AVSS dataset. The best result is highlighted in red.
Model AVSS
$F_{m}$$\uparrow$ $mIoU$$\uparrow$
ResNet-50 3DC .216 17.27
AVS .252 20.18
PCMA .298 24.88
PVT-v2 / Swin AOT .310 25.40
AVS .352 29.77
PCMA .361 31.03
@@ IV SUPERVISION
IoU loss evaluates the similarity between the prediction and GT from a holistic perspective rather than from a single pixel.
@@
V EXPERIMENTS
V-A Datasets and Evaluation Metrics
@@
V-B Implementation Details
TABLE III: Complexity and speed comparison of different AVS algorithms. The result of AVS methods is measured by the default setting. The FPS test is conducted on a 16GB RAM, NVIDIA RTX4060 PC. The best result is highlighted in red. The arrows $\uparrow$ ($\downarrow$) of the metrics indicate the value is higher (lower) the result is better.
Model Params (M)$\downarrow$ FLOPs (G)$\downarrow$ FPS $\uparrow$
ResNet-50 AVS 91.40 169.320 13.83
PCMA 65.68 84.157 32.47
PVT-v2 AVS 101.32 154.137 11.63
PCMA 94.81 107.556 18.47
TABLE IV: Ablation analysis of PCMANet. The best result is highlighted in red. The abbreviations ”GA”, ”CA”, and ”CI” denote the AVGA, QSCA, and CIM modules, respectively. The subscripts ”M”, ”C”, and ”P” represent ”Mask”, ”Continuity”, and ”Progression”, respectively.
Model S4
GA $\text{CA}_{\text{w/o M}}$ $\text{CA}_{\text{w/o C}}$ CA $\text{CI}_{\text{w/o P}}$ CI $F_{m}$$\uparrow$ $mIoU$$\uparrow$
.832 71.16
.846 72.67
.855 74.29
.854 74.68
.853 74.25
.858 75.25
@@ V-B Implementation Details
V-C Comparison with State of the Art Methods
V-C1 Quantitative Comparison
To verify the performance of our method, we conducted a comparative analysis against the state-of-the-art algorithm, AVS [25, 26], and three other related tasks, including sound source localization (SSL), video object segmentation (VOS), and salient object detection (SOD). Specifically, we select several typical algorithms, which are LVS [28] and MSSL [27] for SSL, 3DC [54], SST[55], and AOT [56] for VOS, iGAN [57] and LGVT [58] for SOD. The quantitative results in Table I and II demonstrate that our proposed network outperforms others, using both ResNet-50 and PVT-v2 backbones.
Furthermore, for a comprehensive evaluation, we measured the model sizes of all networks, as presented in Table III. The results highlight that PCMANet achieves superior performance with significantly reduced computational costs and faster processing speeds.
V-C2 Qualitative Comparison
@@
V-D Ablation Analysis
@@
As depicted in Figure IV, we systematically disassembled the modules to assess the impact of each individual component. In the first row, we omitted all modules, and to facilitate a fair comparison, we directly combined the audio and visual features, serving as our baseline. The ”QSCA without mask” condition indicates that we did not perform query selection and instead utilized all the tokens. The ”QSCA without continuity” condition implies that the audio features generated from the QSCA modules were not passed to subsequent QSCA modules but remained constant using initial audio feature $A$. Finally, ”CIM without progression” signifies that the masks calculated by the CIM units were not progressive but treated as independent. The results underscore the effectiveness of the network’s settings and modules, which are well-designed and contribute to its performance.
Figure 8: Mask trends of S4 task during training phase.
Figure 9: Visualization of the inputs, outputs, and attention maps of AVGA module. For the sake of simplicity, the audio wave is omitted in the figure. The x and y coordinate represent $m^{th}$ groups (from $1$ to $8$) and $n^{th}$ stages (from $1$ to $4$), respectively.
V-E Visual Analysis
@@ VI CONCLUSIONS
Sufficient experiments conducted on three benchmark AVS and AVSS datasets demonstrate the superior effectiveness of the proposed method compared to other algorithms. In the future, further investigations could focus on enhancing performance in more complex and challenging tasks, such as AVSS.

### v2

@@ Abstract
Audio and visual signals typically occur simultaneously, and humans possess an innate ability to correlate and synchronize information from these two modalities. Recently, a challenging problem known as Audio-Visual Segmentation (AVS) has emerged, intending to produce segmentation maps for sounding objects within a scene. However, the methods proposed so far have not sufficiently integrated audio and visual information, and the computational costs have been extremely high. Additionally, the outputs of different stages have not been fully utilized. To facilitate this research, we introduce a novel Progressive Confident Masking Attention Network (PMCANet). It leverages attention mechanisms to uncover the intrinsic correlations between audio signals and visual frames. Furthermore, we design an efficient and effective cross-attention module to enhance semantic perception by selecting query tokens. This selection is determined through confidence-driven units based on the network’s multi-stage predictive outputs. Experiments demonstrate that our network outperforms other AVS methods while requiring less computational resources. The code is available at: https://github.com/PrettyPlate/PCMANet.
keywords:
audio-visual segmentation, cross-modal attention mechanism, dynamic network, multi-modality fusion
1 Introduction
@@
As mentioned earlier, the number of unlabeled videos is rapidly increasing. Research on these tasks primarily focuses on semi-supervised learning [18] and unsupervised learning, particularly self-supervised learning [19, 20, 21, 22, 23, 24]. However, due to the lack of patch-level or pixel-level annotation, namely, supervision, these tasks are limited to image-level classification, and their applications are severely constrained. For years, researchers did not pay much attention to supervised learning and pixel-level annotation. Zhou et al. [25, 26] were the first to propose the Audio-Visual Segmentation (AVS) task and Audio-Visual Segmentation with Semantic (AVSS) task, providing benchmark datasets specifically for these tasks. AVS aims to segment the sounding objects in a scene by leveraging both audio and visual modalities, ensuring that the predicted segmentation mask corresponds to the source of the sound. Meanwhile, AVSS extends AVS by assigning semantic labels to segmented objects. There is a similar task called Sound Source Localization (SSL) [20, 18, 21, 27, 22, 23, 28, 29], which aims to locate the regions in the visual scene that are relevant to the audio signals. Although it is not pixel-level, it is annotated at the patch-level. The results are typically represented by heat maps, which are obtained through similarity matrices or Class Activation Mapping (CAM) [30].
@@
2 Related Work
2.1 Audio-Visual Correspondence
@@
2.2 Sound Source Localization
@@
Figure 2: Overview of the pipeline. The network takes the video clip as input and separates it into visual frames and audio spectrograms. The outputs of the audio and visual encoders are denoted as $F_{i}$ and $A$, respectively. The features are first integrated by AVGAs and are further processed by QSCAs to generate guide features $G_{i}$. Meanwhile, the multi-stage outputs, which are aggregated by GFs, will pass through CIMs to generate confidence masks and then be sent to QSCAs for further optimization. The bottom-right corner presents the internal processing logic of the GF module. The deepest GF, which does not receive the feature $H$ as input, is labeled as GF*.
2.3 Audio-Visual Attention
@@
2.4 Audio-Visual Segmentation
Audio-Visual Segmentation (AVS) task aims to generate a pixel-wise segmentation mask of the sounding objects in a scene by leveraging both audio and visual modalities. Unlike conventional segmentation tasks, AVS ensures that the segmented region corresponds to the source of the audio, distinguishing it from non-sounding objects in the scene. Audio-Visual Semantic Segmentation (AVSS) extends AVS by not only segmenting the sounding object but also assigning a semantic category to it.
Zhou et al. [25] introduced the AVS task and later extended it to AVSS [26], establishing benchmark datasets for both tasks. They designed an end-to-end framework for AVS, which adopts the TPAVI module to encode temporal pixel-wise audio-visual interaction. Hao et al. [37] devised an efficient method utilizing bidirectional generation supervision to strengthen correlations between audio-visual modalities. Similarly, Liu et al. [38] introduced the audio-visual semantic correlation (AVSC) module to establish robust associations between audio and visual information. They also introduced a silent object-aware segmentation objective to address limitations in the segmentation supervision of existing AVS datasets. Compared with the one-way feature fusion approach of TPAVI, the proposed method adopts a bidirectional optimization and update strategy for audio and video feature. Compared with the traditional attention computation methods in the aforementioned approaches, the proposed approach introduces a masking and local feature update strategy.
3 Proposed Framework
3.1 Overview Structure of PCMANet
@@
Figure 3: Structure of Audio-Visual Group Attention (AVGA). It integrates audio and visual features using group attention operations. The visual features are divided into multiple groups, with each group fused with audio information through the ”Attn” module. This module first normalizes the input features and then performs fusion using a dot-product operation.
@@
3.2 Audio-Visual Group Attention
With the extracted audio and visual features, it is intuitive to assess whether there exists a correlation between the audio signal and spatial localization. Previous studies[22, 18] introduced an audio-visual attention mechanism by computing the cosine similarity between audio and visual features at the pixel level, and it was found to be effective in tasks such as sound source localization and separation. Drawing inspiration from these findings, we designed a novel module named Audio-Visual Group Attention (AVGA).
@@ 3.2 Audio-Visual Group Attention
3.3 Query-Selected Cross-Attention
Figure 4: Structure of Query-Selected Cross-Attention (QSCA). Right: QSCA takes audio features, visual features, and an external binary mask as inputs and outputs the integrated features. Modal information is exchanged through two multi-head cross-attention blocks, with one employing query-selected attention. Left: Visual queries are selected using gather and scatter operations based on the mask input.
@@ 3.3 Query-Selected Cross-Attention
Furthermore, the visual tokens are redundant, and the computational cost of the attention operation is exceedingly high. To address this challenge, we propose a new query-selected mechanism. Instead of using all tokens, the network selects tokens that lack confidence for segmentation [52]. The criteria for selection are measured by the Confidence-Induced Masking unit (see Section 3.4).
Figure 5: An example of the QSCA and Confidence-Induced Masking (CIM). Dark and light pixel regions represent masked and valid queries, respectively. The mask is generated from the sigmoid output, followed by thresholding and binarization. Progressive masking is achieved through iterative mask multiplication.
@@
where $N^{\prime}=rN$, the ratio $r$ gradually decreases from $1$ to 00. Experiments demonstrate that in the later stages of training, $r$ is lower than $10\%$, indicating that a significant portion of the tokens is masked, resulting in substantial savings in computational resources. For example, in our cases, the last QSCA module takes $28\times 28$ image size input and the channel $C$ is unified to $256$. Thus, we obtain the original MSA FLOPs equals to $520.2M$. If the mask ratio exceeds $90\%$, namely $r<10\%$, the optimized FLOPs of QSCA is less than $144.3M$, which is a third of the original cross-attention computational costs. Figure 5 provides an example of the confidence masks and the selected query.
3.4 Confidence-Induced Masking and Guided Fusion
@@
The Guided Fusion module essentially combines adjacent features ($L$, $H$), similar to FPN [53]. The key distinction lies in its utilization of the output feature $G$ from the QSCA module, which has already incorporated the attention information derived from audio-visual interactions. This modification ensures that the prediction generation process focuses more on the sounding objects. Notably, as shown in Figure 2, the deepest GF module lacks the feature $H$ from the deeper GF module, resulting in only two inputs, $L$ and $G$, and is therefore labeled as GF*.
4 SUPERVISION
@@
IoU loss evaluates the similarity between the prediction and Ground Truth (GT) from a holistic perspective rather than from a single pixel.
@@
5 EXPERIMENTS
5.1 Datasets and Evaluation Metrics
@@
5.2 Implementation Details
@@
5.3 Comparison with State of the Art Methods
5.3.1 Quantitative Comparison
Table 1: Quantitative comparison on AVS dataset. The best result is highlighted in BOLD. The arrow $\uparrow$ of the metrics indicates the value is higher the result is better. ResNet-50, PVT-v2, and Swin refer to the ResNet-50 [39], PVT-v2 [41], and Swin Transformer [55], respectively.
Method Backbone S4 MS3
[…]
LVS ResNet-50 .510 37.94 .330 29.45
MSSL ResNet-18 .663 44.89 .363 26.13
3DC ResNet-34 .759 57.10 .503 36.92
SST ResNet-101 .801 66.29 .572 42.57
TPAVI-R ResNet-50 .848 72.79 .578 47.88
AVSBG ResNet-50 .854 74.13 .568 44.95
PCMA-R (Ours) ResNet-50 .858 75.24 .648 54.76
iGAN Swin-T .778 61.59 .544 42.89
LGVT Swin-T .873 74.94 .593 40.71
TPAVI-T PVT-v2 .879 78.74 .645 54.00
ECMMS Swin-T .886 81.29 .657 59.50
PCMA-T (Ours) PVT-v2 .893 80.10 .708 60.35
Table 2: Quantitative comparison on AVSS dataset. The best result is highlighted in BOLD.
Method Backbone AVSS
[…]
3DC ResNet-34 .216 17.27
TPAVI-R ResNet-50 .252 20.18
PCMA-R ResNet-50 .298 24.88
AOT Swin-T .310 25.40
TPAVI-T PVT-v2 .352 29.77
PCMA-T PVT-v2 .361 31.03
To verify the performance of our method, we conducted a comparative analysis against the state-of-the-art algorithm, TPAVI [25, 26], AVSC [38], AVSBG [37], and three other related tasks, including sound source localization (SSL), video object segmentation (VOS), and salient object detection (SOD). Specifically, we select several typical algorithms, which are LVS [28] and MSSL [27] for SSL, 3DC [56], SST[57], and AOT [58] for VOS, iGAN [59] and LGVT [60] for SOD. The quantitative results in Table 1 and 2 demonstrate that our proposed network outperforms others, using both ResNet and Transformer backbones.
Furthermore, for a comprehensive evaluation, we measured the model sizes of all networks, as presented in Table 3. The results highlight that PCMANet achieves superior performance with significantly reduced computational costs and faster processing speeds.
Table 3: Complexity and speed comparison of different AVS algorithms. The result of AVS methods is measured by the default setting. The FPS test is conducted on a 16GB RAM, NVIDIA RTX4060 PC. The best result is highlighted in BOLD. The arrows $\uparrow$ ($\downarrow$) of the metrics indicate the value is higher (lower) the result is better.
Method Backbone Params (M)$\downarrow$ FLOPs (G)$\downarrow$ FPS $\uparrow$
TPAVI-R ResNet-50 91.40 169.320 13.83
PCMA-R ResNet-50 65.68 84.157 32.47
TPAVI-T PVT-v2 101.32 154.137 11.63
PCMA-T PVT-v2 94.81 107.556 18.47
5.3.2 Qualitative Comparison
@@
5.4 Ablation Analysis
Table 4: Ablation analysis of PCMANet. The best result is highlighted in BOLD. The abbreviations ”GA”, ”CA”, and ”CI” denote the AVGA, QSCA, and CIM modules, respectively. The subscripts ”M”, ”I”, and ”P” represent ”Mask”, ”Interaction”, and ”Progression”, respectively.
Method Module S4
[…]
$\mathcal{M}_{0}$832 71.16
$\mathcal{M}_{1}$846 72.67
$\mathcal{M}_{2}$855 74.29
$\mathcal{M}_{3}$854 74.68
$\mathcal{M}_{4}$853 74.25
$\mathcal{M}_{5}$858 75.25
@@ 5.4 Ablation Analysis
As depicted in Table 4, we systematically disassembled the modules to assess the impact of each individual component. In the first row, we omitted all modules, and to facilitate a fair comparison, we directly combined the audio and visual features, serving as our baseline. The ”QSCA without mask” condition indicates that we did not perform query selection and instead utilized all the tokens. The ”QSCA without interaction” condition implies that the audio features generated from the QSCA modules were not passed to subsequent QSCA modules but remained constant using the initial audio feature $A$. Finally, ”CIM without progression” signifies that the masks calculated by the CIM units were not progressive but treated as independent. The results underscore the effectiveness of the network’s settings and modules, which are well-designed and contribute to its performance.
Figure 8: Ablation study on different fusion strategies. a) Multi-stage fusion and late fusion; b) Interactive fusion and one-way fusion; and c) Masking region update and traditional update.
5.5 Study on Different Fusion Strategies
Moreover, we explore the impact of various fusion strategies to further validate the effectiveness of our model. Specifically, we evaluate three approaches as shown in Figure 8:
  1. a.
    Comparison between multi-stage fusion ($\mathcal{M}_{1}$) and late fusion ($\mathcal{M}_{0}$): This experiment highlights the benefit of applying AVGA fusion across all feature stages, demonstrating the advantages of a multi-stage fusion strategy over late fusion.
  2. b.
    Comparison of interactive audio-visual fusion ($\mathcal{M}_{5}$) and single-way fusion ($\mathcal{M}_{4}$): This analysis emphasizes the importance of bidirectional updates, showing that interactive fusion, where both audio and visual features are iteratively updated, outperforms single-way fusion. Compared to TPAVI (benchmark), which adopts a unidirectional fusion method, the bidirectional interaction approach proposed in this paper offers numerous advantages.
  3. c.
    Comparison of the masking method ($\mathcal{M}_{5}$) and traditional methods ($\mathcal{M}_{2}$): The results indicate that the proposed masking approach improves performance compared to conventional techniques.
These findings collectively validate the design choices in our model and underscore the effectiveness of our proposed fusion strategies.
5.6 Visual Analysis
@@
Figure 9: Mask trends of S4 task during the training phase.
@@
Figure 10: Visualization of the inputs, outputs, and attention maps of AVGA module. For the sake of simplicity, the audio wave is omitted in the figure. The $x$ and $y$ coordinate represent $n^{th}$ groups (from 1 to 8) and $m^{th}$ stages (from 1 to 4), respectively.
@@ 6 CONCLUSIONS
Sufficient experiments conducted on three benchmark AVS and AVSS datasets demonstrate the superior effectiveness of the proposed method compared to other algorithms. The computational efficiency and robust performance of PCMANet make it highly applicable to real-world scenarios, such as field or industrial video surveillance and edge computing systems. In the future, further investigations could focus on enhancing performance in more complex and challenging tasks, such as AVSS.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported by the China Postdoctoral Science Foundation under Grant 2023M741952.
Appendix A More Visualization
Figure 11: Visualization of the intermediate features before and after the fusion of audio and visual information.
Fig. 11 illustrates an additional visualization of the intermediate features before and after the fusion. The middle column presents the visual features. The top row displays features derived from visual frames alone (vision-only), while the bottom row shows the fused features after integrating audio signals. The visual features without fusion tend to highlight all objects within the scene indiscriminately. In contrast, the fused features are more focused on regions corresponding to objects associated with the audio signal (e.g., the dog in the first example and the drum in the second). This observation indicates that audio signals act as a guiding mechanism, directing the visual features to concentrate on relevant regions.